The reciprocal of half a number increased by half the reciprocal of a number is one half. Find the number...please :)
Answer by amoresroy(361)

The reciprocal of half a number increased by half the reciprocal of a number is one half. Find the number...please :)
Hi
Let x = the number
x/2 = half the number
2/x = the reciprocal of half the number

1/2x= half the reciprocal of the number
Here's the equation:
2/x + 1/2x = 1/2
Multiply both sides of equation by 2x to cancel denominator
4 + 1 = x
x = 5
The number is 5
-----------------------
Regards,
Roy Amores